Find the general solution of yuy-xux=2
1
Expert's answer
2021-09-28T10:03:51-0400

"\frac{dx}{-x}=\frac{dy}{y}=\frac{du}{2}"

  1. First equation

"\frac{dx}{-x}=\frac{dy}{y}\\
\int \frac{dx}{-x}=\int \frac{dy}{y}\\
ln|xy|=ln]C|\\
xy=C;"

2 second equation

"\frac{dy}{y}=\frac{du}{2}\\
\int \frac{dy}{y}=\int \frac{du}{2}\\
ln|y|+C=\frac{u}{2}\\
u-ln(y^2)=C"

So we have two integrals of the charactestic system:

"xy=C_1;"

"u-ln(y^2)=C_2"

Therefore general solution has form

"F\left( xy, u-ln(y^2) \right)=0"

or "u=ln(y^2)+H\left( x\cdot y \right)"

where H(t)- any differentable function